• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 00:30:17 on May 07, 2010

    Entry number 321297

    keyword=Compton Cavity Work Summary

    1. Accelerator vacuum crew replaced the leaky valve VBV1P02. (9:00-13:00) 2. Optimized PPLN profile and alignment (13:00-14:00) 3. Installed a HeNe laser and coincided the red beam with green beam to diagnose future misalignment problem (13:00-16:00) 4. Tried to install Leysop FOI, with a new backet from machine shop, even after the alignment and rotation angle tweak, it didn't give good transmission. (15:00-16:30) 5. Replaced and aligned the cavity mirrors and obtained 60+% optical coupling. Put together the bellows and closed up the cavity. Cavity is being pumped with turbo pump, maintained the resonance during the pumping. (17:00- 23:30)

    Tomorrow we will do a leak check, if it passes the leak check we will do bake out for 24 hours.

    Mohammed & Abdurahim